OpenStudy (anonymous):

Given the domain values { -2, 0, 2 }, find the range values for the function y = 5x + 1.

OpenStudy (goformit100):

put the value of x in the given equation

OpenStudy (goformit100):

try it and reply

OpenStudy (anonymous):

5(-2)+1=-9 5(0)+1=1 5(2)+1=11